OpenStudy (anonymous):

Log Properties Question: When solving a logarithmic equation, say one of the terms on the LHS is log, and the only term on the RHS is 1. What gives me the right to drop the log from the LHS and replace the RHS with 10?

OpenStudy (anonymous):

As in the following example: \[logx + \log(x+3) = 1\] \[\log(x^2+3x) = 1\] \[x^2+3x = 10\]

OpenStudy (anonymous):

if the base is 10 :)